a.Sales of $ 3000000 are subject to an accrued warranty cost of 8​%. The accrued warranty payable at the beginning of the year was $ 40000​, and warranty payments for the year totaled $ 54000.

b.On March ​1, Jenkins Electronics signed a $ 20000 note payable that requires annual payments of $ 4000 plus 7​% interest on the unpaid balance beginning March ​1, 2019.

c.Panorama​, ​Inc., a chain of discount​ stores, ordered $ 130000 worth of wireless speakers and related products. With its​ order, Panorama​, ​Inc., sent a check for $ 130000 in​ advance, and Jenkins shipped $ 45000 of the goods. Jenkins will ship the remainder of the goods on April ​3, 2018.

d.Jenkins​'s March payroll of $ 270000 is subject to employee withheld income tax of $ 30000 and FICA tax of 7.65​%. On March 31​, Jenkins pays employees their​ take-home pay and accrues all tax amounts

Report these items on Jenkin ​Electronics' balance sheet at March 31 2018.

Current liabilities:

Accrued warranty payable

Interest payable

Current portion of long-term note payable

Unearned sales revenue

Employee withheld income tax payable

FICA tax payable

Total current liabilities

Long-term liabilities:

Note payable
Answer #1

1              CURRENT LIABILITIES:   

A ACCRUED WARRANTY PAYABLE A/C        $2,26,000

B             INTEREST PAYBALE          $117

C CURRENT PORTION OF LONG TERN NOTE PAYABLE          $4,000

D             UNEARNED SALES REVENUE       $85,000

E EMPLOYEE INCOME TAX PAYABLE            $30,000

F              FICA TAX PAYABLE           $20,655

                TOTAL CURRENT LIABILITIES        $3,65,772

                               

2              LONG TERM LIABILITIES

                NOTE PAYABLE $16,000

WORKING NOTES:

a.                           

SALES A/C           $30,00,000         

WARRANTY EXPENSES A/C         $2,40,000           

WARRANTY PAYABLE A/C                            

OPENING            $40,000               

ADD: ADDITIONS              $2,40,000           

LESS: PAYMENTS              $-54,000             

CLOSING              $2,26,000           

                               

b.                           

LOAN AMOUNT                $20,000               

INTEREST ACCURED ON 31ST MARCH 2019           =20000*7%*1/12             $117

FIRST ANNUAL PAYMENT ON 1ST MARCH, 2019 = CURRENT PORTION OF LONG TERN NOTE PAYABLE = $4,000

NOTE PAYABLE IN LONG TERM = 20000-4000 = $16000  

                               

c.                            

ORDER VALUE $1,30,000           

GOODS SHIPPED              $45,000               

UNEARNED SALES REVENUE       =130000-45000 =$85000

IF IT IS ASSUMED THAT SALES IS RECORDED AT THE TIME OF SHIPMENT                

                               

                               

d.                           

TOTAL (GROSS) SALARY $2,70,000           

INCOME TAX      $30,000               

FICA TAX PAYABLE (7.65%)          $20,655               

NET SALARY        $2,19,345           

                               

TAX ACCRUALS = INCOME TAX $30,000